how do you solve (by elimination)
4x - y = -3
5x-2y=18

Answer by ewatrrr(24785)

 
Hi
solving (by elimination)
4x - y = -3
5x-2y=18
-8x + 2y = 6 |multiplying 1st EQ by -2 to eliminate y
5x - 2y = 18
-3x = 24
x = -8 and y = -29 (y = 4*-8 + 3)
CHECKING our Answer***
5x - 2y = 18
-40 + 58 = 18